We are the only Bubble Tea vendor on the IU Bloomington campus. Right now we are charging a price of $4 per bubble tea to all our customers. However, we have noticed that student inverse demand P1(Q1) differs from faculty/staff inverse demand P2(Q2), as seen by the given weekly demand curves:

Students: P1(Q1) = 10 – 0.01Q1

Faculty/Staff: P2(Q2) = 5 – 0.01Q2

In an effort to increase our profits we are thinking about customizing prices to the two different customer groups (3rd degree price discrimination). I am inclined to lower prices to students as they are always looking for a bargain! But please review both groups' prices and see if you spot any opportunities for enhancing profits. In your review, keep in mind that our marginal costs per bubble tea served are $1.
Answer #1

Inverse demand curve for students : P1=10-0.01Q1

Now, total revenue TR1=P1*Q1 = 10Q1-0.01Q12

Then, marginal revenue MR1=dTR1/dQ1 = 10-0.02Q1

For equilibrium, MR1 = MC

or, 10-0.02Q1=1

or, 0.02Q1=9

or, Q1=450 units

and P1=10-(0.01*450) = $5.5

Similarly, inverse demand curve for faculty : P2=5-0.01Q2

Now, total revenue TR2=P2*Q2 = 5Q2-0.01Q22

Then, marginal revenue MR2=dTR2/dQ2 = 5-0.02Q2

For equilibrium, MR2 = MC

or, 5-0.02Q2=1

or, 0.02Q2=4

or, Q2=200 units

and P2=5-(0.01*200) = $3

Then profit in price discrimination = P1Q1+P2Q2-{MC*(Q1+Q2)} = (5.5*450)+(3*200)-{1*(450+200)} = 2475+600-650 = $2425